# On the backward stability of s-step GMRES

### v1

@@ Abstract
Communication, i.e., data movement, is a critical bottleneck for the performance of classical Krylov subspace method solvers on modern computer architectures. Variants of these methods which avoid communication have been introduced, which, while equivalent in exact arithmetic, can be unstable in finite precision. In this work, we address the backward stability of $s$-step GMRES, also known as communication-avoiding GMRES. We present a framework for simplifying the analysis of $s$-step GMRES, which includes standard GMRES ($s=1$) as a special case, by isolating the effects of rounding errors in the QR factorization and the solution of the least squares problem. Using this framework, we analyze $s$-step GMRES with popular block orthogonalization methods: block modified Gram–Schmidt and reorthogonalized block classical Gram–Schmidt algorithms.
An example illustrates the resulting instability of $s$-step GMRES when paired with the classical $s$-step Arnoldi process and shows the limitations of popular strategies for resolving this instability. To address this issue, we propose a modified Arnoldi process that allows for much larger block size $s$ while maintaining satisfactory accuracy, as confirmed by our numerical experiments.
@@ 1 Introduction
In each step, the $s$-step GMRES algorithm generates a block Krylov basis consisting of $s$ vectors formed as follows:
@@
Considering rounding errors, the backward stability of the standard GMRES algorithms with Householder and modified Gram–Schmidt orthogonalization (MGS) has been investigated in [12, 19]. Recently, [5] introduced a comprehensive framework to simplify the rounding error analysis of GMRES algorithms. Building on this framework, we analyze the rounding error of the $s$-step GMRES algorithm to illustrate how the condition number of the basis affects the backward error in this study. Our analysis is performed in a way so that it is possible identify the sources of errors, and clarifying the impact of different orthogonalization methods on the backward error of the computed $s$-step GMRES results and broadening the applicability of the framework introduced by [5]. In addition, our analysis formally shows that the backward error of $s$-step GMRES is influenced by the condition number of the basis. From this analysis, we further propose a modified Arnoldi process to allow for the use of a substantially larger block size $s$, as confirmed by numerical experiments, increasing the cost of the orthogonalization by a factor of 2.
The reminder of this paper is organized as follows. In Section 2, we give an overview of the $s$-step Arnoldi and GMRES algorithms. In Section 3, we present our abstract framework for the analysis of the $s$-step GMRES method. In Section 4, we invoke our abstract framework to analyze the backward error of the $s$-step GMRES algorithm with different commonly-used block orthogonalization methods. We discuss the implications of the theoretical results and comment on stopping criteria in Section 5. Then we propose a modified $s$-step Arnoldi process in Section 6 to improve the backward stability of the $s$-step GMRES. Numerical experiments are presented in Section 7 which compare the $s$-step GMRES algorithm with the modified and classical $s$-step Arnoldi processes.
@@
2 $s$-step Arnoldi and GMRES algorithms
@@
The standard GMRES algorithm, namely $s$-step GMRES algorithm with $s=1$, employs the Arnoldi process to build an orthonormal basis $V_{1:i+1}$ of the Krylov subspace $\mathcal{K}_{i+1}(A,r)$ with $r=b-Ax^{(0)}$, satisfying
@@
with $W_{1:i}=AV_{1:i}$ and the upper triangular matrix $R_{1:i+1}$, after $i$ iterations, and then update $x^{(i)}\leftarrow x^{(0)}+V_{1:i}y^{(i)}$ if converged, where $y^{(i)}=\operatorname*{\arg\min}_{y}\bigl{\|}\beta e_{1}-H_{1:i+1,1:i}y\bigr{\rVert}$ with $\beta=\bigl{\|}r\bigr{\rVert}$, $e_{1}$ denoting the first column of the $(i+1)$-by-$(i+1)$ identity matrix, and $H_{1:i+1,1:i}=R_{1:i+1,2:i+1}$.
@@ 3 An abstract framework for backward stability of $s$-step GMRES
In the recent work [5], the authors propose a framework to study the backward stability of what they call “modular GMRES” as follows:
@@
To assess the backward stability of GMRES with various orthogonalization methods, it is necessary to evaluate the rounding errors in Step 2 for the chosen orthogonalization method within this framework. Note that Step 2 in modular GMRES comprises orthogonalization, Givens QR factorization, and solving a triangular system. Therefore, an analysis of all three parts is required when modifying the orthogonalization method used in GMRES, which is less convenient and unnecessary, and further causes ambiguity regarding the requirements of the orthogonalization method.
In this section, we extend the above framework to $Z_{1:(i+1)s}=M_{R}^{-1}B_{1:(i+1)s}$ focusing on the backward stability of the $s$-step GMRES method. This includes Algorithm 2 with the classical $s$-step Arnoldi process (Algorithm 1), and the standard GMRES algorithm, which is essentially a specific case of $s$-step GMRES with $s=1$ and $p_{0}(A)=I$, hence $B_{1:(i+1)s}=V_{1:(i+1)s}$. Furthermore, we delve into the block orthogonalization, Givens QR decomposition, and triangular system solving steps, which are actually performed in the algorithm, as opposed to analyzing $y^{(i+1)}=\operatorname*{\arg\min}_{y}\bigl{\|}M_{L}^{-1}b-W_{1:(i+1)s}y\bigr{\rVert}$. From this analysis, we will present the assumptions only relevant to the block orthogonalization process itself. Consequently, we only need to consider the rounding errors in the orthogonalization process when modifying the orthogonalization method utilized in the GMRES algorithm. We note that the approach here thus represents an improved version of the modular GMRES framework in [5] for the standard (non $s$-step) GMRES case.
@@
    Solve the least squares problem $y^{(i+1)}=\operatorname*{\arg\min}_{y}\bigg{\|}R_{1:(i+1)s+1}\begin{bmatrix}1\\ -y\end{bmatrix}\bigg{\rVert}$, which is through solving the triangular system $T_{1:(i+1)s}y^{(i+1)}=\beta G_{1,1:(i+1)s}^{\top}$ to obtain $y^{(i+1)}$, where $T_{1:(i+1)s+1,1:(i+1)s}$ is $R$-factor of the QR factorization by Givens rotations, i.e.,
@@
Taking rounding errors into account, we further assume that each line satisfies the following, to be described below. In the remainder of this subsection, $\varepsilon_{*}$, $\delta_{*}$, and $\omega_{*}$ denote constants in the rounding error analysis of different operations, which helps to identify the source of errors in the final results. We use $\delta$ to indicate constants associated with the left and right preconditioners, and use $\omega$ to denote constants relevant to the block orthogonalization process. Constants that are independent of these two parts are denoted by $\varepsilon$, which can be considered as $\varepsilon_{*}\in(0,1)$, as ensured by standard rounding error analysis; see, e.g., [13].
@@
$\widehat{Z}_{1:(i+1)s}$ $=M_{R}^{-1}\hat{B}_{1:(i+1)s}+\Delta Z_{1:(i+1)s},\quad\bigl{\|}\Delta Z_{j}\bigr{\rVert}\leq\delta_{M_{R}^{-1}}\bigl{\|}M_{R}^{-1}\bigr{\rVert}\bigl{\|}\hat{B}_{j}\bigr{\rVert},$ (3)
@@
for any $j\in\{1,2,\dotsc,(i+1)s\}$. Let $\Delta W_{1:(i+1)s}=M_{L}^{-1}\Delta C_{1:(i+1)s}+\Delta D_{1:(i+1)s}$. $\widehat{W}_{1:(i+1)s}$ further satisfies
@@
where $\widehat{Z}_{1:(i+1)s}=\tilde{Z}_{1:(i+1)s}D_{1:(i+1)s}$ with a diagonal positive definite matrix $D_{1:(i+1)s}$.
@@
Then we assume the backward stability, which is satisfied by most commonly-used block orthogonalization schemes summarized in [10], of the block orthogonalization method used in Step 2; i.e., for any $\phi>0$,
@@
for any $j\in\{2,\dotsc,(i+1)s+1\}$, which implies
@@
In addition, we define $\omega_{is+j}$ to describe the loss of orthogonality of the $Q$-factor, i.e., for any $j\in\{1,2,\dotsc,s\}$,
$\begin{split}&\bigl{\|}\bar{V}_{1:is+j}^{\top}\bar{V}_{1:is+j}-I\bigr{\rVert}_{\mathsf{F}}\leq\omega_{is+j},\\ &\bigl{\|}\bar{V}_{1:is+j}\bigr{\rVert}\leq\sqrt{\bigl{\|}I\bigr{\rVert}+\bigl{\|}\bar{V}_{1:is+j}^{\top}\bar{V}_{1:is+j}-I\bigr{\rVert}}\leq\sqrt{1+\omega_{is+j}}.\end{split}$ (12)
@@
About the least squares problem, [5] measures $\min_{y}{\bigl{\|}\widehat{r}-\widehat{W}_{1:p}y\bigr{\rVert}}$. It is essential to note that we avoid accessing $\min_{y}{\bigl{\|}\widehat{r}-\widehat{W}_{1:p}y\bigr{\rVert}}$, due to the substantial impact of the block orthogonalization process on this residual. In this context, we aim to distinguish the errors introduced by Steps 2 and 3 separately. Therefore, our attention is on (15), which is performed in Step 3 and remains unaffected by the block orthogonalization.
@@
From standard rounding error analysis and [13, Equation (3.12)], we have $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{mv}=O(\bm{u})$. In [5, Equation (A.7)], the authors have given $\varepsilon_{ls}=O(\bm{u})$, which is from Givens QR factorization and the triangular solver.
@@
To bound $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$, we first need to establish the relationship between the result and the assumptions (13), (14), which amounts to bounding
$\sigma_{\min}\bigg{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\bigg{)}\quad\text{and}\quad\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$
by $\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$ and $\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}$. We summarize the result in the following lemma.
@@
Before proving Lemma 1, we give the following lemma which will be used in the following proofs.
@@
Given $X\in\mathbb{R}^{n\times m}$. Assume that $X+\Delta X=QU$, where $Q\in\mathbb{R}^{n\times m}$ satisfies $\bigl{\|}Q^{\top}Q-I\bigr{\rVert}_{\mathsf{F}}<1$. Then
@@
From (9), if $\bar{V}_{1:p+1}$ is exact orthonormal, we have
@@
$\begin{split}\bar{V}_{1:p+1}&\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\\ &=\bar{V}_{1:p}\begin{bmatrix}\phi\widehat{R}_{1:p,1}&\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}\end{bmatrix}+\bar{V}_{p+1}\widehat{R}_{p+1,1:p+1}D_{1:p}^{-1}.\end{split}$
From (9), we obtain
@@
with $\Delta F=\Delta E_{1}(\phi)+\Delta E_{1:p+1}(\phi)D_{1:p}^{-1}-\bar{V}_{p+1}\widehat{R}_{p+1,1:p+1}D_{1:p}^{-1}$ satisfying
$\bigl{\|}\Delta F\bigr{\rVert}_{\mathsf{F}}\leq\omega_{\mathrm{qr}}\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$
which also implies, by (13) and Lemma 2,
$\begin{split}&\big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p,1}&\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}]\end{bmatrix}\big{\rVert}\\ &\qquad\leq\frac{1+\omega_{p}}{1-\omega_{p}}\bigl{(}(1+\omega_{\mathrm{qr}})\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}\\ &\qquad\leq 3\bigl{(}(1+\omega_{\mathrm{qr}})\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}\end{split}$ (30)
and
@@
Together with
@@
It remains to prove (20) and (21). From Lemma 2 and

### v2

@@ Abstract
Communication, i.e., data movement, is a critical bottleneck for the performance of classical Krylov subspace method solvers on modern computer architectures. Variants of these methods which avoid communication have been introduced, which, while equivalent in exact arithmetic, can be unstable in finite precision. In this work, we address the backward stability of $s$-step GMRES, also known as communication-avoiding GMRES. Compared to the “modular framework” proposed in [A. Buttari, N. J. Higham, T. Mary, & B. Vieublé. Preprint in 2024.], we present an improved framework for simplifying the analysis of $s$-step GMRES, which includes standard GMRES ($s=1$) as a special case, by isolating the effects of rounding errors in the QR factorization and the solution of the least squares problem. The key advantage of this new framework is that it is evident how the orthogonalization method affects the backward error, and it is not necessary to re-evaluate anything other than the orthogonalization itself when modifying the orthogonalization used in GMRES. Using this framework, we analyze $s$-step GMRES with popular block orthogonalization methods: block modified Gram–Schmidt and reorthogonalized block classical Gram–Schmidt algorithms.
An example illustrates the resulting instability of $s$-step GMRES when paired with the classical $s$-step Arnoldi process and shows the limitations of popular strategies for resolving this instability. To address this issue, we propose a modified $s$-step Arnoldi process that allows for much larger block size $s$ while maintaining satisfactory accuracy, as confirmed by our numerical experiments.
@@ 1 Introduction
In each iteration, the $s$-step GMRES algorithm generates a block Krylov basis consisting of $s$ vectors formed as follows:
@@
Considering rounding errors, the backward stability of the standard GMRES algorithms with Householder and modified Gram–Schmidt orthogonalization (MGS) has been investigated in [12, 19]. Recently, [5] introduced a comprehensive framework to simplify the rounding error analysis of GMRES algorithms. Building on this framework, we analyze the rounding error of the $s$-step GMRES algorithm to illustrate how the condition number of the basis affects the backward error in this study. Our analysis is presented with the goals of making it possible to easily identify the sources of errors, clarifying the impact of different orthogonalization methods on the backward error of the computed $s$-step GMRES results, and broadening the applicability of the framework introduced by [5]. In addition, our analysis formally shows that the backward error of $s$-step GMRES is influenced by the condition number of the basis. From this analysis, we further propose a modified Arnoldi process to allow for the use of a substantially larger block size $s$, as confirmed by numerical experiments, increasing the cost of the orthogonalization by a factor of 2.
The remainder of this paper is organized as follows. In Section 2, we give an overview of the $s$-step Arnoldi and GMRES algorithms. In Section 3, we present our abstract framework for the analysis of the $s$-step GMRES method. In Section 4, we invoke our abstract framework to analyze the backward error of the $s$-step GMRES algorithm with different commonly-used block orthogonalization methods. We discuss the implications of the theoretical results and comment on stopping criteria in Section 5. Then we propose a modified $s$-step Arnoldi process in Section 6 to improve the backward stability of the $s$-step GMRES. Numerical experiments are presented in Section 7 which compare the $s$-step GMRES algorithm with the modified and classical $s$-step Arnoldi processes.
@@
2 The $s$-step Arnoldi and GMRES algorithms
@@
The standard GMRES algorithm, namely the $s$-step GMRES algorithm with $s=1$, utilizes the Arnoldi process to build an orthonormal basis $V_{1:i+1}$ for the Krylov subspace $\mathcal{K}_{i+1}(A,r)$, where $r=b-Ax^{(0)}$. This process is expressed by
@@
where $W_{1:i}=AV_{1:i}$ and $R_{1:i+1}$ is an upper triangular matrix. After $i$ iterations, if convergence is attained, the solution is updated via $x^{(i)}\leftarrow x^{(0)}+V_{1:i}y^{(i)}$ with $y^{(i)}=\operatorname*{\arg\min}_{y}\bigl{\|}\beta e_{1}-H_{1:i+1,1:i}y\bigr{\rVert}$. Here, $\beta=\bigl{\|}r\bigr{\rVert}$, $e_{1}$ represents the first column of the identity matrix of size $(i+1)\times(i+1)$, and $H_{1:i+1,1:i}=R_{1:i+1,2:i+1}$.
@@ 3 An abstract framework for backward stability of $s$-step GMRES
In the recent work [5], the authors propose a framework to study the backward stability of what they call the “modular GMRES” algorithm, which is as follows:
@@
This modular GMRES framework is capable of capturing a wide range of GMRES variants, including the standard GMRES algorithm with various orthogonalization methods. By analyzing the rounding errors in each step, [5] demonstrates that this modular GMRES algorithm is backward stable under mild assumptions. Among these three steps, Steps 1 and 3 are typically straightforward to analyze using standard rounding error analysis. To assess the backward stability of GMRES with various orthogonalization methods, it is necessary to evaluate the rounding errors in Step 2 for the chosen orthogonalization method within this framework. Note that there are implicit algorithmic choices in Step 2 based on various orthogonalization methods. Therefore, Step 2 needs to be thoroughly re-evaluated when modifying the orthogonalization method used in GMRES, which is inconvenient (and unnecessary), and further causes ambiguity regarding the requirements of the orthogonalization method. It is possible to expand the analysis from [5] to the $s$-step GMRES method by substituting the orthonormal basis $V_{1:(i+1)s}$ with a general basis $B_{1:(i+1)s}$. However, due to the above-mentioned deficiencies, we will not fully follow the analysis presented in [5].
In this section, we develop an improved framework for $Z_{1:(i+1)s}=M_{R}^{-1}B_{1:(i+1)s}$ focusing on the backward stability of the $s$-step GMRES method. This includes Algorithm 2 with the classical $s$-step Arnoldi process (Algorithm 1), and the standard GMRES algorithm, which is essentially a specific case of $s$-step GMRES with $s=1$ and $p_{0}(A)=I$, hence $B_{1:(i+1)s}=V_{1:(i+1)s}$. Furthermore, we delve into the block orthogonalization, Givens QR decomposition, and triangular system solving steps, which are actually performed in the algorithm, as opposed to analyzing $y^{(i+1)}=\operatorname*{\arg\min}_{y}\bigl{\|}M_{L}^{-1}b-W_{1:(i+1)s}y\bigr{\rVert}$. From this analysis, we will present the assumptions only relevant to the block orthogonalization process itself. Consequently, we only need to consider the rounding errors in the orthogonalization process when modifying the orthogonalization method utilized in the GMRES algorithm. We note that the approach here thus represents an improved version of the modular GMRES framework in [5] for the standard (non $s$-step) GMRES case.
@@
    Solve the least squares problem $y^{(i+1)}=\operatorname*{\arg\min}_{y}\bigg{\|}R_{1:(i+1)s+1}\begin{bmatrix}1\\ -y\end{bmatrix}\bigg{\rVert}$, which is through solving the triangular system $T_{1:(i+1)s}y^{(i+1)}=\beta G_{1,1:(i+1)s}^{\top}$ to obtain $y^{(i+1)}$, where $T_{1:(i+1)s+1,1:(i+1)s}$ is the $R$-factor of the QR factorization by Givens rotations, i.e.,
@@
Taking rounding errors into account, we further assume that each line satisfies the following, to be described below. In the remainder of this subsection, $\varepsilon_{*}$, $\delta_{*}$, and $\omega_{*}$ denote constants in the rounding error analysis of different operations, which helps to identify the source of errors in the final results. We use $\delta$ to indicate constants associated with the left and right preconditioners, and use $\omega$ to denote constants relevant to the block orthogonalization process. Constants that are independent of these two parts are denoted by $\varepsilon$, which can be considered as $\varepsilon_{*}\in(0,1)$, as ensured by standard rounding error analysis with the assumption $n\bm{u}\ll 1$; see, e.g., [13].
@@
$\widehat{Z}_{1:(i+1)s}$ $=M_{R}^{-1}\widehat{B}_{1:(i+1)s}+\Delta Z_{1:(i+1)s},\quad\bigl{\|}\Delta Z_{j}\bigr{\rVert}\leq\delta_{M_{R}^{-1}}\bigl{\|}M_{R}^{-1}\bigr{\rVert}\bigl{\|}\widehat{B}_{j}\bigr{\rVert},$ (3)
@@
for any $j\in\{1,2,\dotsc,(i+1)s\}$. Let $\Delta W_{1:(i+1)s}=M_{L}^{-1}\Delta C_{1:(i+1)s}+\Delta D_{1:(i+1)s}$. Then we obtain
@@
where $\widehat{Z}_{1:(i+1)s}=\tilde{Z}_{1:(i+1)s}D_{1:(i+1)s}$ with any invertible diagonal matrix $D_{1:(i+1)s}$.
@@
In GMRES, the orthogonalization method employed in Step 2 can be viewed as being applied to the matrix $\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:(i+1)s}\end{bmatrix}$ for any positive scalar $\phi$. We assume the backward stability of the block orthogonalization method used in Step 2; according to [10], most commonly-used block orthogonalization schemes satisfy this property, i.e., for any $\phi>0$,
@@
for any $j\in\{2,\dotsc,(i+1)s+1\}$. Furthermore, (10) implies
@@
In addition, we define $\omega_{is+j}$ to describe the loss of orthogonality of $\bar{V}_{1:is+j}$, i.e., for any $j\in\{1,2,\dotsc,s\}$,
$\begin{split}\bigl{\|}\bar{V}_{1:is+j}^{\top}\bar{V}_{1:is+j}-I\bigr{\rVert}_{\mathsf{F}}&\leq\omega_{is+j},\\ \bigl{\|}\bar{V}_{1:is+j}\bigr{\rVert}&=\sqrt{\bigl{\|}\bar{V}_{1:is+j}^{\top}\bar{V}_{1:is+j}\bigr{\rVert}}\leq\sqrt{\bigl{\|}I\bigr{\rVert}+\bigl{\|}\bar{V}_{1:is+j}^{\top}\bar{V}_{1:is+j}-I\bigr{\rVert}}\\ &\leq\sqrt{1+\omega_{is+j}}.\end{split}$ (12)
@@
Regarding the least squares problem, [5] measures $\min_{y}{\bigl{\|}\widehat{r}-\widehat{W}_{1:p}y\bigr{\rVert}}$, which involves Steps 2 and 3 of the algorithm. Thus, the substantial impact of the block orthogonalization process, i.e., Step 2, will be unclear when assessing $\min_{y}{\bigl{\|}\widehat{r}-\widehat{W}_{1:p}y\bigr{\rVert}}$ directly. In this context, we avoid directly accessing $\min_{y}{\bigl{\|}\widehat{r}-\widehat{W}_{1:p}y\bigr{\rVert}}$ and aim to distinguish the errors introduced by Steps 2 and 3 separately. Therefore, our attention is on (15), which is performed in Step 3 and remains unaffected by the block orthogonalization.
@@
We summarize the quantities $\epsilon_{*}$, $\delta_{*}$, and $\omega_{*}$, coming from different operations in Table 1.
Table 1: Notation of rounding error analysis for Steps 1, 2, 3, and 4.
Notation Sources Details
$\delta_{M_{R}^{-1}}$ Applying $M_{R}^{-1}$ to $B_{1:(i+1)s}$ (3)
$\varepsilon_{AZ}$ Computing $AZ_{1:(i+1)s}$ (4)
$\delta_{M_{L}^{-1}AZ}$ Applying $M_{L}^{-1}$ to $AZ_{1:(i+1)s}$ (4)
$\delta_{M_{L}^{-1}b}$ Applying $M_{L}^{-1}$ to $b$ (7)
$\omega_{\mathrm{qr}}$ Orthogonalizing $\begin{bmatrix}\phi r&W_{1:(i+1)s}\end{bmatrix}$ (9)
$\omega_{is+j}$ Loss of orthogonality of $V_{1:is+j}$ (12)
$\omega_{\kappa(r,W)}$ The upper bound on $1/\kappa\bigl{(}\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\bigr{)}$ to show if $r$ lies in the range of $W_{1:p}$ (13)
$\omega_{\kappa(W)}$ The lower bound on $1/\kappa(\widehat{W}_{1:p}D_{1:p}^{-1})$ to show if the columns of $W_{1:p}$ are linearly independent (14)
$\varepsilon_{ls}$ Solving the least squares problem $\min_{y}\bigg{\|}R_{1:p+1}\begin{bmatrix}1\\ -y\end{bmatrix}\bigg{\rVert}$ (15)
$\varepsilon_{mv}$ Computing $Z_{1:p}y^{(i^{*})}$ (17)
@@
From standard rounding error analysis and [13, Equation (3.12)], we have $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{mv}=O(\bm{u})$. In [5, Equation (A.7)], the authors have given $\varepsilon_{ls}=O(\bm{u})$, which comes from applying Givens QR factorization to the upper Hessenberg matrix $\hat{R}_{1:p+1,2:p+1}$ and solving the triangular system as depicted in Step 3.
@@
According to [19, Theorem 2.4], the residual of the least squares problem in Step 3, i.e., $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$, can be bounded by $\sigma_{\min}\bigg{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\bigg{)}$ and $\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$. Therefore, establishing a bound for the residual $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$ requires determining the relationship between the result and the assumptions (13) and (14), which amounts to bounding
$\sigma_{\min}\bigg{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\bigg{)}$
by $\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$ and bounding
$\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$
by $\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}$. We summarize the result in the following lemma.
@@
Notice that $\hat{R}_{1:p+1}$ is obtained as the $R$-factor by employing the block QR factorization on $\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:(i+1)s}\end{bmatrix}$ as indicated in (9). Therefore, before proving Lemma 1, we present the following lemma to estimate the largest and smallest singular values of the computed $R$ factor.
@@
Given $X\in\mathbb{R}^{n\times m}$, assume that $X+\Delta X=QU$, where $Q\in\mathbb{R}^{n\times m}$ satisfies $\bigl{\|}Q^{\top}Q-I\bigr{\rVert}_{\mathsf{F}}<1$. Then
@@
Notice that $\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}\end{bmatrix}$ is the computed $R$-factor of $\begin{bmatrix}\phi r,\widehat{W}_{1:p}\end{bmatrix}$ from (9). We then multiply two sides of  (9) by $\begin{bmatrix}1&0\\ 0&D^{-1}_{1:p}\end{bmatrix}$ to obtain
$\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}+\Delta\tilde{E}_{1:p+1}(\phi)=\bar{V}_{1:p+1}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix},$ (29)
where $\Delta\tilde{E}_{1:p+1}(\phi):=\Delta E_{1:p+1}(\phi)\begin{bmatrix}1&0\\ 0&D^{-1}_{1:p}\end{bmatrix}$ satisfies (11).
From (29), if $\bar{V}_{1:p+1}$ is exactly orthonormal, we have
@@
$\begin{split}\bar{V}_{1:p+1}&\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\\ &=\bar{V}_{1:p}\begin{bmatrix}\phi\widehat{R}_{1:p,1}&\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}\end{bmatrix}+\bar{V}_{p+1}\widehat{R}_{p+1,1:p+1}D_{1:p}^{-1}.\end{split}$ (30)
By substituting (30) into the right-hand side of (29), we derive
@@
with $\Delta F=\Delta\tilde{E}_{1:p+1}(\phi)-\bar{V}_{p+1}\widehat{R}_{p+1,1:p+1}D_{1:p}^{-1}$ satisfying, from (11) and the definition (26) of $\omega_{H}$,
$\bigl{\|}\Delta F\bigr{\rVert}_{\mathsf{F}}\leq\omega_{\mathrm{qr}}\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}.$
In (31), notice that $\begin{bmatrix}\phi\widehat{R}_{1:p,1}&\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}\end{bmatrix}$ is the $R$-factor of $[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]+\Delta F$. Using (13) and the perturbation results of QR factorization, i.e., Lemma 2, we have the bound on the largest singular value,
$\begin{split}&\big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p,1}&\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}]\end{bmatrix}\big{\rVert}\\ &\qquad\leq\frac{1+\omega_{p}}{1-\omega_{p}}\bigl{(}(1+\omega_{\mathrm{qr}})\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}\\ &\qquad\leq 3\bigl{(}(1+\omega_{\mathrm{qr}})\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)},\end{split}$ (32)
and the smallest one,
@@
By substituting these two bounds and (26) into the right-hand side of the two inequalities
@@
It remains to prove (20) and (21), which is similar to the proof of (18) and (19). From Lemma 2 and